### v2

# A Constructive Definition of Space via Dynamical Evolution and Observational Acts

Abstract
We propose a constructive and dynamical redefinition of spatial structure, grounded in the interplay between mechanical evolution and observational acts. Rather than presupposing space as a static background, we interpret space as an emergent entity that arises through observational acts. Using the framework of pre-topologies, measurable structures, and the GNS construction, we analyze how the choice of observables and the system’s time evolution dynamically determine the topological and measure-theoretic features of space. This approach highlights the observer-dependent and context-sensitive nature of spatial concepts in both classical and quantum domains.
Nature and Environment, Faculty of Liberal Arts, The Open University of Japan, Chiba 261-8586, Japan
1 Introduction
In mathematics, the concept of “space” appears in a wide variety of forms—topological spaces, metric spaces, measure spaces, linear spaces, Hilbert spaces, and more. Despite their diversity, these structures are all called “spaces”, often without a clearly articulated reason for their common designation. While the mathematical operations applied within these spaces are rigorously defined, the notion of space itself is often left to intuitive understanding, heavily dependent on context. Thus, the term “space” functions not so much as a rigorously defined mathematical object, but rather as a conceptual apparatus—one that resides outside formal mathematics and is anchored in the intuition and contextual judgment of the mathematician.
This conceptual ambiguity not only signals a lack of foundational clarity, but also allows implicit philosophical assumptions to infiltrate mathematical reasoning—often without scrutiny.
Even in rigorous disciplines such as functional analysis or algebraic geometry—fields that emphasize internal consistency—the notion of “space” often functions more as an intuitive scaffold than as a strictly defined object.
In light of these issues, we ask: What is space—not as a physical or metaphysical object, but as a mathematical construct shaped by the observer’s methodology and purpose? The mathematical concept of space, we argue, is deeply rooted in the mathematician’s observational acts and constructive motives. It is perhaps more accurate to see it as a reflection of the mathematician’s spatial intuition than as a strictly defined object.
In what follows, we approach this question constructively: we reinterpret spatial structures not as static entities, but as dynamical constructs shaped by time evolution and the act of observation.
By exploring this question, we aim to shed light on the internal structure of the spatial concept in mathematics and reveal that even abstract mathematical spaces are often constrained within the classical framework of space that we intuitively understand. This line of thought aligns with Kato’s Leibnizian relational view of space[1], as well as modern topos-theoretic perspectives, in which space is not a pre-existing container but a structure that emerges as a result of constructive activity.
Mathematics is an indispensable tool for physics. Since Newton, and through the development of relativity and quantum theory, physics has evolved beyond the classical intuition of space and time. Meanwhile, mathematics—particularly through foundational studies originating from problems like the heat equation—has taken a different path. Although the two disciplines occasionally interact in contexts such as gauge theory or general relativity, they largely follow separate trajectories. In this context, it is not evident that modern mathematics functions as an appropriate tool for modern physics.
For instance, quantum field theory and path integral formulations still lack fully rigorous mathematical definitions. Moreover, when fundamental questions regarding the nature of spacetime are raised, traditional spatial concepts inherited from classical physics may no longer be valid—particularly in light of Bohr’s complementarity and the breakdown of classical realism in quantum contexts.
With this in mind, the present paper proposes a constructive redefinition of space that is grounded in physical dynamics and observational acts. We begin with a reinterpretation of topological concepts from the perspective of dynamical systems, introducing the notion of reachability domains and pre-topologies. We then consider measurable structures as projections determined by observation, and finally, we examine the role of non-commutative observables in quantum theory. Through these investigations, we show how spatial structures arise dynamically and contextually, depending on the system’s evolution and the observer’s choices.
This perspective bears resemblance to several prior approaches, yet differs in crucial ways that highlight its novelty.
A related line of inquiry can be found in Connes’ noncommutative geometry [5], which extends the Gel’fand–Naimark duality—originally linking commutative $C^{*}$-algebras and topological spaces—to the noncommutative setting. In Connes’ framework, geometry is encoded algebraically via $C^{*}$-algebras, leading to the notion of “quantum spaces.” However, this formulation lacks an explicit account of the dynamical and observational aspects that are central to our approach. In contrast, we argue that it is precisely the time evolution of a system and the observer’s choice of observable that jointly give rise to spatial structures. The topology of space is dynamically determined by the system’s evolution, while the measure structure emerges from the act of observation. In quantum theory, this leads naturally to the context-dependence of mutually incompatible observables—i.e., the impossibility of simultaneous measurement. While Connes’ formalism assumes a fixed topological and measure-theoretic backdrop within a $C^{*}$-dynamical system, our proposal allows these structures to vary depending on the dynamical system and the observer’s choices [^1] While the essential contribution of Connes’ approach is the introduction of the Dirac operator to define geometric structures, his framework does not explicitly address the dynamical and observation-dependent aspects emphasized here, and is thus more static in character..
The groundbreaking aspect of Connes’approach lies in the observation that, once the algebra of “functions” on space becomes noncommutative, the very notion of a point loses its conventional meaning. In this sense, noncommutative geometry deals with what is often called a point-less geometry. By contrast, in our constructive framework, the set of observables under consideration is always taken to be a mutually commuting set within a chosen context, so that within this restricted set the assignment of values is still meaningful. In this sense, the notion of a “point” retains a partial operational meaning.
A similar spirit is found in Rovelli’s relational interpretation of quantum mechanics [9, 11], which also emphasizes the constructive role of the observer. However, Rovelli’s interpretation relies on Everett’s “relative state” formulation [12], treating only relational facts as ontologically meaningful. In this view, any physical system may serve as an observer for another. Similar ideas have been extended to classical mechanics in [13]. Nonetheless, these approaches emphasize relational events rather than time evolution as a structuring principle. In our formulation, by contrast, temporal dynamics play a central role in defining pre-topological and topological structures. Furthermore, whereas Rovelli emphasizes the relational nature of observers, we emphasize that the choice of observable by the observer is what directly shapes the emerging spatial structure. This distinction marks a significant departure from prior approaches and underscores the fundamentally constructive stance taken throughout this work.  [^2] An earlier version of this work was originally published in Japanese as [3].
As another line of discussion, following Born’s proposal of the reciprocity principle[14], there has been recent interest in quantum gravity scenarios that take seriously the possibility that momentum space itself may be curved[15][16]. Such approaches often regard phase space, rather than spacetime, as the fundamental arena, with the dynamical nature of the observer influencing the reconstruction of space[17]. This perspective shares common ground with our view that space should be constructed from dynamics and observation. However, from our standpoint, momentum is regarded merely as one possible observable, and in this sense our formulation may be considered more general.
2 Pre-Topological Spaces and Reachability Domains: A Theory of Space Without Observers
In this section, we revisit the concept of topological space from the standpoint of pre-topology, using the notion of reachability domains derived from the time evolution of dynamical systems. We will demonstrate how such reachability structures naturally induce a pre-topological structure on a set. Subsequently, by analyzing these reachability domains, we provide a dynamical interpretation of topological spaces themselves. From this perspective, a topological space can be understood as a kind of equilibrium state—where opposing temporal behaviors achieve a balance—within a dynamical system.
2.1 Dynamical Systems
A dynamical system is a theoretical framework for describing physical quantities that evolve over time. Let $x(t)$ denote the state of a system at time $t$. If we denote the initial state at time $t=0$ as $x_{0}$, then the time evolution of the system after $t$ units of time can be represented as
$x(t)=U(x_{0},t).$ (2.1)
Here, $U$ is a function representing the dynamical evolution. Although this notation may resemble the unitary evolution operator in quantum theory, it should be noted that $U$ is not assumed to be unitary in general. That is, we are not restricting ourselves to Hamiltonian systems [^3] Note that the only primitives assumed here are a set and its dynamics; no metric or pre-existing topological structure is presupposed.
Importantly, the term “dynamical system” is used here in a broad sense—it refers not only to Hamiltonian mechanics, but to any system that exhibits time-dependent evolution.
By definition, the evolution function satisfies:
$U(x_{0},0)=x_{0},$ (2.2)
which simply asserts that the state remains unchanged at time zero.
The trajectory of a point $x(t)$ under such a system, given an initial state $x_{0}$, a dynamical evolution $U$, and a time horizon $T$, is defined by the set
$D_{U}(x_{0},T)=\{U(x_{0},t)|0\leq t\leq T\}.$ (2.3)
This trajectory depends on the initial condition, the dynamics of the system, and the duration of evolution. If the system evolves in discrete time, then $T\in\mathbb{Z}$, whereas for continuous systems, we take $T\in\mathbb{R}$.
This immediately raises the question: what kinds of structures can serve as “time”? At the very least, it is conceivable that time could be modeled not only by the real or integer numbers but also by more general algebraic structures such as monoids, groups, or partially ordered sets (posets).
Although one could consider more exotic scenarios—such as systems with multiple time parameters or higher-dimensional time—we restrict ourselves here to the simplest case: a single, linear notion of time444 Generalization to the case where time is modeled as a poset is technically straightforward, and would naturally connect to Lieb-Robinson-type causal structures in many-body quantum systems, suggesting further links to Lieb’s thermodynamic formulation[18]. A detailed treatment is beyond the scope of the present work..
When time is modeled as a monoid, repeated composition of time evolutions becomes well-defined and naturally interpretable. In such a case, time admits a binary operation (typically denoted by composition or addition) and a unit element, enabling the successive application of the evolution operator.
Let us consider a dynamical evolution operator
$U(t):X\to X,$ (2.4)
where $x\in X$ evolves under $U$. If time is structured as a monoid, then the following properties hold:
$U(s)\circ U(t)=U(s\circ t),$ (2.5)
$U(e)=id_{x},$ (2.6)
where $e$ is the identity element of the monoid, and $\text{id}_{x}$ is the identity map on $X$. In this case, since inverse elements are not assumed to exist, the “arrow of time” is irreversible.
If time is modeled instead by a group, the existence of inverse elements allows for time-reversal symmetry: for every $t$, there exists t-t such that U(t)U(-t) is inverse of $U(t)$.
Alternatively, if time is represented by a partially ordered set (poset), the time evolution respects a partial order structure. This interpretation naturally encodes causal relationships—events evolve in a way that preserves causal ordering.
This discussion can be summarized in the following slogans:
  • Monoid: Linear repetition of time steps $t_{1},t_{2},t_{3}$
  • Group: Symmetry under time reversal $t\longleftrightarrow-t$
  • Poset: Branching and causal ordering through partial order
In most physical theories, time is treated as a totally ordered set, and we will adopt that assumption here as well. However, it is important to note that in subsequent sections—especially when discussing reachability domains, closure, and continuity—the nature of time (whether totally ordered or structured as a general monoid) will influence how these topological notions are defined and interpreted.
The nature of the space $X$, to which the evolving quantity $x(t)\in X$ belongs, can vary widely depending on the physical or mathematical context. If $X$ is taken to be real space and the dynamical system is governed by a Hamiltonian, we are in the setting of classical mechanics. If $X$ represents a thermodynamic phase space and the dynamics follow Onsager’s theory, the framework corresponds to nonequilibrium thermodynamics.
More generally, $X$ may be an abstract set or a topological space. In the context of quantum mechanics, for instance, $X$ can be interpreted as a space of operators. Thus, the framework of dynamical systems is flexible enough to encompass a wide variety of physical and mathematical theories by appropriate choices of the underlying space $X$ and the dynamics $U$.
2.2 Reachability Domains
We now introduce the concept of reachability, which plays a central role in our constructive interpretation of space.
Given a dynamical system $U$ and a time horizon $T$, the reachability domain of an initial region $X_{0}\subset X$ is defined as the union of the trajectories emanating from all initial points in $X_{0}$ over the time interval $[0,T]$.
Formally, we define:
$\bar{X}_{0}^{T}(U)\equiv D_{U}(X_{0},T)\equiv\{D_{U}(x_{0},T)|x_{0}\in X_{0}\}$ (2.7)
$=\{U(x_{0},t)|x_{0}\in X_{0},0\leq t\leq T\}=\mathrm{cl}_{U}^{T}(X_{0}).$ (2.8)
Here,
$D_{U}(x_{0},T):=\{\,U(x_{0},t)\mid 0\leq t\leq T\,\}.$ (2.9)
We use the overline notation $\bar{X}_{0}^{T}$ as a shorthand for $\mathrm{cl}_{U}^{T}(X_{0})$, to suggest a closure operation, anticipating the later identification of this object with a kind of dynamical closure in the context of pre-topological structures. Here $\mathrm{cl}_{U}^{T}$ plays the role of a closure operator; if it satisfies idempotency, the induced structure is a topology, otherwise a pre-topology.
Conceptually, this construction describes how a “bundle of particles” initially distributed over $X_{0}$ evolves over time under the dynamics $U$. As we shall see, such reachability domains form the foundation for inducing spatial structures dynamically—space emerges from the collective evolution of matter or information.
2.3 Example
Let us now consider a simple illustrative example. Take the set
$X=\{1,2,3,4,5\}.$ (2.10)
Define a dynamical system $U$ such that the state shifts one position to the right at each unit time step:
$U(x,1)=x+1.$ (2.11)
We impose periodic boundary conditions so that
$U(5,1)=1.$ (2.12)
Let the initial region be
$X_{0}=\{1,2,3\}.$ (2.13)
Then the reachability domain after one time step is:
$\bar{X}_{0}^{T=1}(U)=\{1,2,3,4\}.$ (2.14)
This example illustrates how the structure of the reachability domain evolves under the dynamics. A key observation is that when a subset coincides with its reachability domain, it can be regarded as a closed set in the pre-topological sense. In the next section, we develop this interpretation more formally, defining closed and open sets within a pre-topological framework based on dynamical behavior.
2.4 Dynamical Interpretation of Pre-Topology
A pre-topological space on a set $X$, denoted as $(\{X_{0}\subset X\},U,T)$, is defined by equipping XXwith a dynamical system $U$ and a time horizon $T$, and by examining how subsets $X_{0}\subset X$ evolve over time.
Specifically, the structure of the pre-topology is determined by how the reachability domains of initial regions $X_{0}$ behave under the dynamics up to time $T$.
[…]
A key feature of this perspective is that the structure of the pre-topology depends on the dynamics. By modifying the dynamical system $U$, the pre-topological structure on $X$ also changes. Likewise, altering the time horizon $T$ affects the reachability domains and thus reshapes the structure of the pre-topological space.
In this way, we begin to see a direct link between dynamics and spatial structure: space is no longer a fixed backdrop but a system-dependent, emergent structure shaped by temporal evolution.
2.5 Dynamical Interpretation of Closed Sets
The notion of a closed set can be naturally reinterpreted through a dynamical lens. In this context, a subset $X_{0}\subset X$ is considered closed if, under the evolution of the dynamical system, the region remains invariant—that is, no part of it “escapes” under time evolution. Formally, this condition is expressed as:
$\bar{X}_{0}^{T}(U)=X_{0}.$ (2.15)
In this formulation, a closed set corresponds to a dynamically stable region—one that is preserved under the time evolution dictated by the system $U$. Thus, the classical topological definition of a closed set gains a physical interpretation: it is a region from which the dynamical flow cannot exit over the time interval $[0,T]$.
As a simple example, consider the set
$X=\{1,2,3,4\},$ (2.16)
$X_{0}=\{2,4\},$ (2.17)
and define the dynamics as
$U(x,1)=x+2,$ (2.18)
with periodic boundary conditions. Then, after one time step, we have:
$\overline{X}_{0}^{T}=\{2,4\}=X_{0}$ (2.19)
which satisfies the condition for $X_{0}$ to be a closed set under this dynamical system.
2.6 Dynamical Interpretation of Interiors and Open Sets
An initial region $X_{0}\subset X$ may either remain invariant under time evolution—that is, satisfy the closedness condition $\bar{X}_{0}^{T}(U)=X_{0}$ —or expand over time, so that $X_{0}\subset\bar{X}_{0}^{T}(U)$. Similarly, the complement of the initial region, denoted $X_{0}^{c}$, may also either remain invariant $\overline{X_{0}^{c}}^{T}=X_{0}^{c}$ or expand: $X_{0}^{c}\subset\overline{X_{0}^{c}}^{T}(U)$.
When the complement region $X_{0}^{c}$ remains unchanged under the evolution, that is,
$\overline{X_{0}^{c}}^{T}=X_{0}^{c},$ (2.20)
we define the original set $X_{0}$ to be an open set with respect to the dynamics.
The interior of a set $X_{0}$, denoted $X_{0}^{\circ}$, is then defined by taking the complement of the closure of its complement:
$X_{0}^{\circ}\equiv\left(\overline{X_{0}^{c}}^{T}\right)^{c}.$ (2.21)
This definition aligns with classical topology but is now grounded in a dynamical interpretation: a set is open if the flow does not “intrude” into its complement over the time interval considered.
As a concrete example, let us revisit the case from the previous section involving a closed set. Suppose
$X=\{1,2,3,4\},$ (2.22)
$X_{0}=\{2,4\}.$ (2.23)
Then its complement is
$X_{0}^{c}=\{1,3\},$ (2.24)
and the closure of this complement under the dynamics is
$\overline{X^{c}}_{0}^{T}=\{1,3\}$ (2.25)
which remains unchanged. Therefore, $X_{0}$ satisfies the condition for being an open set.
The interior of $X_{0}$, according to the definition, is:
$X_{0}^{o}=\left(\overline{X_{0}^{c}}^{T}\right)^{c}=\{2,4\}.$ (2.26)
Let us now consider another example. Let
$X=\{1,2,3,4,5\},$ (2.27)
$X_{0}=\{1,2,3\},$ (2.28)
and define the time evolution as:
$U(x,1)=x+1,$ (2.29)
with periodic boundary conditions $U(5,1)=1$.
Then the complement is
$X_{0}^{c}=\{4,5\},$ (2.30)
and its closure under one time step is:
$\overline{X_{0}^{c}}^{T}=\{1,4,5\}.$ (2.31)
Since this differs from $X_{0}^{c}$, $X_{0}$ is not an open set under this dynamical system.
Its interior is given by:
$X_{0}^{o}=\left(\overline{X_{0}^{c}}^{T}\right)^{c}=\{2,3\}.$ (2.32)
2.7 Dynamical Interpretation of Topological Spaces
From the preceding discussion, we can now reinterpret the notion of a topological space from a dynamical perspective. Given a set $X$, a dynamical system $U$, and a time horizon $T$, we define a topological structure by collecting those subsets $X_{0}\subset X$ whose reachability domains under $U$ up to time $T$ coincide with the original set:
$\bar{X}_{0}^{T}(U)=X_{0}.$ (2.33)
That is, a subset is considered closed if it remains invariant under the system’s time evolution. Taking complements leads to a corresponding definition of open sets. In both cases, the defining property is that the evolution does not “leak out” of the subset or into its complement, respectively.
This condition reflects a kind of dynamical equilibrium: the initial region and its complement are in a stable, mutually restrictive relationship. From this viewpoint, a topological space appears not as a static background, but as a set of regions stabilized through the temporal dynamics of the system.
In the next section, we could naturally proceed to redefine notions such as continuity and homeomorphism between pre-topological spaces in dynamical terms. However, since the focus of this paper is on the constructive origin of space itself, we omit such technical developments here and refer interested readers to [3] for further elaboration.
3 Measurable Structures and the Act of Distinction in Observation
In this section, we explore how measurable structures can be understood as arising from projective structures associated with observation. That is, a measurable structure is not something inherently present in space, but rather reflects the observer’s decision about what distinctions are to be made within that space.
From this viewpoint, observation is the act of instituting distinctions: it is the selection of criteria by which one portion of a space is separated from another. The measurable structure, then, encodes the observer’s choice of which features or properties are to be considered relevant and discernible.
Thus, we propose that measurement is not merely the extraction of information from a pre-existing structure, but a constructive act that determines the structure itself—by selecting what distinctions to recognize, the observer effectively determines the measurable geometry of the space.
3.1 Measurable Spaces
Let $X$ be a given set. Suppose we assign a property to each element $x\in X$ by means of a binary function:
$p(x)=\begin{cases}1&\mathrm{if\ }x\mathrm{\ satisfies\ the}\ \mathrm{property,}\\ 0&\mathrm{otherwise}.\end{cases}$ (3.1)
This can be viewed as a projection of the elements of $X$ onto the Boolean values $\{0,1\}$, indicating whether or not each element satisfies the given property.
Based on this projection, we can define a subset of $X$ consisting of all elements that satisfy the property:
$A=p^{-1}(1),$ (3.2)
$A^{c}=p^{-1}(0).$ (3.3)
In this way, a single property $p$ induces a partition of $X$ into two regions: those that satisfy the property and those that do not.
To refine the distinctions within the space, we may consider multiple properties, denoted $p_{1},p_{2},\dots,p_{n}$. Each property $p_{i}$ gives rise to a corresponding measurable set:
$A_{i}=p_{i}^{-1}(1).$ (3.4)
Each $A_{i}$ is the preimage of $1$ under the projection $p_{i}$, representing the set of points in $X$ for which the $i$-th property holds.
Each set $A_{i}=p_{i}^{-1}(1)$, corresponding to a particular property $p_{i}$, can stand in various logical relationships to other sets $A_{j}=p_{j}^{-1}(1)$ defined by other properties $p_{j}$. If we consider, for example, just two properties $p_{1}$ and $p_{2}$, then we can form combinations of the corresponding sets to reflect various observational queries about the system.
These combinations form a collection $\mathcal{B}$ of subsets of $X$, given by:
[…]
where $\not${O} is the empty set. These subsets encode all the possible distinctions that can be made by combining the two properties.
This collection $\mathcal{B}$ is called a $\sigma$-algebra(sigma-algebra), meaning that it is closed under the operations of complement, countable union, and countable intersection.
A pair $(X,\mathcal{B}$), where $\mathcal{B}$ is $\sigma$-algebra of subsets of $X$, is called a measurable space [^5] The point here is not the textbook definition of a $\sigma$-algebra per se, but the interpretation of this structure as arising from distinctions instituted through acts of observation.. In our context, this $\sigma$-algebra reflects the observer’s selection of which distinctions (i.e., which properties) are relevant and observable.
As a concrete example, consider the finite set:
$X=\{1,2,3,4,5\}.$ (3.6)
Let us define two properties to distinguish elements of $X$:
  • $p_{e}(x)=1$ if $x$ is even, and 0 otherwise;
  • $p_{o}(x)=1$ if $x$ is prime number, and 0 otherwise.
Then, the corresponding subsets defined by these properties are:
$A_{e}=p_{e}^{-1}(1)=\{2,4\},\ A_{e}^{c}=p_{e}^{-1}(0)=\{1,3,5\},$ (3.7)
$A_{p}=p_{p}^{-1}(1)=\{2,3,5\},\ A_{p}^{c}=p_{p}^{-1}(0)=\{1,4\}.$ (3.8)
From these, we can generate a $\sigma$-algebra $\mathcal{B}$ by forming all Boolean combinations of $A_{e}$ and $A_{p}$. The resulting collection includes:
[…]
$A_{e}^{c}\cup A_{p}^{c},A_{e}^{c}\cap A_{p}^{c},A_{e}^{c}\cap A_{p},A_{e}\cap A_{p}^{c},X\}$
[…]
$\{1,3,4,5\},\{1\},\{3,5\},\{4\},X\}.$
Each of these subsets corresponds to a particular distinction made by combining the properties of evenness and primality. For instance, the set of elements that are even but not prime is given by:
$A_{e}\cap A_{p}^{c}=\{4\}.$ (3.10)
This illustrates how observational distinctions—based on multiple binary properties—can generate a $\sigma$-algebra, and hence a measurable structure, on a given space.
3.2 Using Topological Spaces as Structures of Distinction
When considering how to implement distinctions within a set, it is natural to turn to the framework of topological spaces. In the context of our earlier discussion, recall that the reachability domain of a region $X_{0}$ was equal to $X_{0}$ itself. Thus, selecting a subset $X_{0}$ corresponds to selecting a measurable set $A$ as in the previous section.
In this way, the collection of such subsets corresponds to what is known as the Borel $\sigma$-algebra—that is, the $\sigma$-algebra generated by open sets in a topological space. Under this interpretation, a division of the set $X$ into distinct regions implies that the physical quantities within each region do not leak into the other regions under the time evolution. The distinctions are preserved dynamically.
This observation leads immediately to a natural generalization: even in the pre-topological setting—where the reachability domain does not necessarily coincide with the initial region—we can still construct a measurable space $(X,\mathcal{B})$. In this case, the elements of $\mathcal{B}$ are given by reachability domains, which encode the dynamically accessible regions from various initial conditions.
In the next section, we introduce the notion of measure, which will serve as the foundation for defining states—both in classical systems and in quantum theory. The assignment of a measure reflects the observer’s expectation or weighting over the distinctions represented by measurable sets.
3.3 Introducing Measures: The Foundation of the Concept of State
Thus far, we have treated space as something that emerges dynamically: a structure shaped by the time evolution of a system, represented through reachability domains. However, all of this structure has been kinematic—it tells us what regions are dynamically possible, but says nothing about how likely each possibility is to occur.
In this view, a topological space encodes which trajectories are possible under the system’s dynamics. But to describe the actual behavior of a system—or more precisely, its state—we must specify how these possibilities are weighted. This is where the notion of measure enters: it assigns a numerical value to each region, quantifying how much the system is expected to occupy or pass through that region.
$\mu:\mathcal{B}\to[0,1],\ \mu(X_{0})\in\mathbb{R}_{\geq 0}.$ (3.11)
When interpreted probabilistically, the entire space satisfies
$\mu(X)=1,$ (3.12)
and the measure defines a state of the system. The key insight is that the state is not an intrinsic object, but arises from assigning possibility weights to dynamically accessible regions. In this sense, the measure is not something added on top of space—it is the quantitative expression of the system’s relationship to its space.
Now, let us take an element $X_{0}\in\mathcal{B}$ from $\sigma$-algebra of measurable sets, and denote it by $x$. An observable can be understood as a function $f(x)$ defined over such measurable subsets—i.e., a real-valued function that assigns values to outcomes of observation.
The expectation value of an observable f(x^)f($\hat${x}) , given a measure $\mu$, is written as:
$\omega(f(\hat{x}))\equiv\langle f(\hat{x})\rangle_{\rho}=\int f(x)d\mu(x)=\int f(x)\rho(x)dx$ (3.13)
Here $\rho$ is the probability density corresponding to the measure $\mu$, and $\omega$ is the state, in the sense of functional analysis or *-algebra theory. The variable $\hat{x}$ represents a random variable, or observable, whose expectation is determined by this measure-theoretic structure.
The triple $(X,\mathcal{B},\mu)$ is thus called a measure space. Within this framework, we can now clearly state:
A state is a map that assigns a distribution of possibility (i.e., a measure) to a structure of observation (i.e., the $\sigma$-algebra $\mathcal{B}$).
In this way, the state is not a static label, but the quantified outcome of choosing how to distinguish—how to carve up the space into observable regions. The measure reflects both the dynamical potentiality of space (through selection of observables).
3.4 Noncommutative Space and Quantum Theory
The construction of a measure space, as described above, naturally extends to cases where the observables are noncommutative—that is, when they cannot be simultaneously measured or diagonalized. The technical details of this extension are deferred to Appendix A, where we describe how a Hilbert space arises via the GNS (Gelfand–Naimark–Segal) construction once a state and a dynamical algebra are given.
From the perspective of quantum systems, the essential difference between classical and quantum theories lies in the commutativity of observables. In classical theory, all observables commute, and a single global measure space suffices to describe the system. In contrast, in quantum theory, observables may not commute, and only subsets of commuting observables can be simultaneously assigned measurable structure.
When an observer selects a particular observable $\hat{x}$, only those observables that commute with $\hat{x}$ can meaningfully define a simultaneous measurement context. The corresponding measure space $)(X(\hat{x}),\mathcal{B}(\hat{x}),\mu(\hat{x}$)) emerges as the projection of the system onto this selected observable context. In other words, the “space” that becomes visible to the observer is the one constructed by the act of choosing $\hat{x}$.
Thus, even in the quantum case, spatial structure remains fundamentally tied to the observer’s act of selection. What appears as “space” is not a fixed background but a context-dependent construct—one shaped by the algebraic properties of observables and the epistemic stance of the observer.
The constructive nature of spatial structure—and its noncommutative generalization—are elaborated in detail in Appendix A.
3.5 Constructive Dynamics and Measurement in Spin Precession
As a concrete example, we consider the precession of a spin in an external magnetic field. We demonstrate how the dynamical evolution of the spin gives rise to a reachability domain, and how the observer’s choice of observable affects not only the associated measurable structure but also the reachability domain itself [^6] While the analysis of spin precession is standard, our discussion differs in that it interprets the resulting trajectories as reachability domains whose structure depends on the choice of observable, thereby providing a concrete example of space emerging from dynamics and measurement..
When spin is present, its interaction with an external magnetic field is described by the Hamiltonian
$\hat{H}_{I}=-\hat{\mu}_{i}B^{i}$ (3.14)
where $\mu_{i}$ denotes the magnetic moment of the particle, and $B^{i}$ represents the external magnetic field.
For an electron, the spin-induced magnetic moment is given by
$\hat{\mu}_{i}=-g\mu_{B}\frac{\hat{\psi}_{i}}{\hbar}$ (3.15)
where $g\approx 2$ is the $g$-factor for a free electron, and $\mu_{B}=\frac{e\hbar}{2m_{e}}$ is the Bohr magneton.
The operator $\hat{\psi}_{i}$ is defined by
$\hat{\psi}_{i}=\frac{\hbar\sigma_{i}}{2}$ (3.16)
where $\sigma_{i}$ are the Pauli matrices, satisfying the commutation and anticommutation relations
$\sigma_{i},\ \sigma_{j}]=2i\epsilon_{ijk}\sigma_{k},\ \{\sigma_{i},\sigma_{j}\}=2\delta_{ij}.$ (3.17)
From this, the corresponding commutation and anticommutation relations for $\hat{\psi}_{i}$ are
$\hat{\psi}_{i},\ \hat{\psi}_{j}]=i\hbar\epsilon_{ijk}\hat{\psi}_{k},\ \{\hat{\psi}_{i},\hat{\psi}_{j}\}=\frac{\hbar^{2}}{2}\delta_{ij}.$
In the literature, the spin operator $\hat{\psi}_{i}$ is often denoted as $\hat{S}_{i}$.
Now, we consider a situation in which the particle is either extremely massive or subject to a strong external magnetic field, such that the free Hamiltonian $\hat{H}_{0}$ can be neglected. That is, we approximate
$\hat{H}=\hat{H}_{0}+\hat{H}_{I}\approx-\hat{\mu}_{i}B^{i}=-g\frac{e\hbar}{2m_{e}}\frac{\hat{\psi}_{i}}{\hbar}B_{i}=-g\frac{e\hbar B_{i}}{2^{2}m_{e}}\sigma_{i}.$ (3.18)
We now consider the time evolution of the spin observable $\hat{\psi}_{i}$. In the Heisenberg picture, the time-dependent operator is given by
$\hat{\psi}_{i}(t)=\hat{U}^{\dagger}(t)\hat{\psi}_{i}\hat{U}(t),$ (3.19)
where the unitary evolution operator is
$\hat{U}(t)=e^{-i\frac{\hat{H}t}{\hbar}}=\exp\left(ig\frac{et}{2m_{e}\hbar}\hat{\psi}_{i}B_{i}\right)=\exp\left(i\frac{geB_{i}}{4m_{e}}t\sigma_{i}\right).$ (3.20)
We define
$\tilde{B}^{i}\equiv\frac{geB_{i}}{4m_{e}},\ \tilde{B}\equiv|\tilde{\bm{B}}|,\ n^{i}\equiv\frac{\tilde{B}^{i}}{\tilde{B}}$, (3.21)
so that the unitary operator becomes
$\hat{U}(t)=\exp\left(i\tilde{B}^{i}t\sigma_{i}\right)=I\cos\left(\tilde{B}t\right)+in^{i}\sigma_{i}\sin\left(\tilde{B}t\right).$ (3.22)
Since the unitary action of Pauli matrices corresponds to a rotation by twice the angle in vector space, we have
$\hat{\psi}_{i}(t)=\hat{U}^{\dagger}(t)\hat{\psi}_{i}\hat{U}(t)=R^{ij}(2\tilde{B}t)\hat{\psi}_{j}$ (3.23)
where $R^{ij}(\theta)$ is the $\mathrm{SO}(3)$ rotation matrix by angle $\theta$ around axis $\vec{n}$.
The infinitesimal time evolution is then governed by the Heisenberg equation of motion:
$i\hbar\frac{d\hat{\psi}_{i}}{dt}=[\hat{H},\hat{\psi}_{i}]=[-g\frac{e\hbar}{2m_{e}}\frac{\hat{\psi}_{j}}{\hbar}B^{j},\hat{\psi}_{i}].$ (3.24)
Using the commutation relations, this becomes
$\hat{\psi}_{j}B^{j},\hat{\psi}_{i}]=i\hbar\epsilon_{jik}B^{j}\hat{\psi}_{k},$ (3.25)
and we obtain
$\frac{d\hat{\psi}_{i}}{dt}=g\frac{e}{2m_{e}}\epsilon_{ijk}B_{j}\hat{\psi}_{k}.$ (3.26)
This is the precession equation for the spin operator in a magnetic field.
[…]
$\hat{\psi}_{\pm}=\frac{1}{\sqrt{2}}(\hat{\psi}_{1}\pm i\hat{\psi}_{2})$ (3.27)
which satisfy $\hat{\psi}_{-}|0\rangle=0$ where $|0\rangle$ denotes the spin-down state.
The time-evolved state vector is given by
$|\Psi(t)\rangle=\hat{U}(t)|0\rangle=\cos(\tilde{B}t)|0\rangle+i\sin(\tilde{B}t)\left(\frac{2}{\hbar}n^{i}\hat{\psi}_{i}\right)|0\rangle.$ (3.28)
This can be further decomposed using the basis $|0\rangle$, $|1\rangle$ as
$|\Psi(t)\rangle=$ $\left(\cos(\tilde{B}t)+i\sin(\tilde{B}t)\frac{2}{\hbar}n^{i}\langle 0|\hat{\psi}_{i}|0\rangle\right)|0\rangle$ (3.29)
$+\left(i\sin(\tilde{B}t)\frac{2}{\hbar}n^{i}\langle 1|\hat{\psi}_{i}|0\rangle\right)|1\rangle.$
Since the spin operators are related to the Pauli matrices via
$\hat{\psi}_{i}=\frac{\hbar\sigma_{i}}{2},$ (3.30)
we adopt the standard representation of the Pauli matrices:
$\sigma_{1}=\left(\begin{array}[]{cc}0&1\\ 1&0\end{array}\right),\ \sigma_{2}=\left(\begin{array}[]{cc}0&-i\\ i&0\end{array}\right),\ \sigma_{3}=\left(\begin{array}[]{cc}1&0\\ 0&-1\end{array}\right).$ (3.31)
Under this representation, the time-evolved spin state can be written as a two-component spinor on the Bloch sphere:
$|\Psi(t)\rangle=\left(\cos\left(\tilde{B}t\right)-i\sin(\tilde{B}t)n^{3}\right)|0\rangle+i\sin\left(\tilde{B}t\right)(n^{1}-in^{2})|1\rangle=\left(\begin{array}[]{c}\cos\left(\tilde{B}t\right)-in^{3}\sin(\tilde{B}t)\\ i(n^{1}-in^{2})\sin\left(\tilde{B}t\right)\end{array}\right).$ (3.32)
One can verify that this state remains normalized:
$\langle\Psi(t)|\Psi(t)\rangle=\cos^{2}\left(\tilde{B}t\right)+\sin^{2}\left(\tilde{B}t\right)=1.$ (3.33)
Thus, the trajectory of the spin state under time evolution describes a circular motion on the Bloch sphere, centered around the axis defined by the unit vector $\vec{n}$.
Since the trajectory traced by the state vector is a closed circle on the Bloch sphere, the reachability domain for times $t>\frac{2\pi}{\tilde{B}}$ becomes the circle $S^{1}$ around the axis defined by the vector $\vec{n}$.
Now consider the initial state $|0\rangle$, which satisfies
$\hat{\psi}_{3}|0\rangle=-\frac{\hbar}{2}|0\rangle.$ (3.34)
Using the time evolution operator, we find
$\hat{U}(t)\hat{\psi}_{3}\hat{U}(t)^{-1}\hat{U}(t)|0\rangle=-\frac{\hbar}{2}\hat{U}(t)|0\rangle,$ (3.35)
which implies
$\hat{\psi}_{3}(t)|\Psi(t)\rangle=-\frac{\hbar}{2}|\Psi(t)\rangle.$ (3.36)
This result means the following: if the observable is co-rotated with the state—that is, we choose $\hat{\psi}_{3}(t)$ as the observable at time tt—then the system remains in an eigenstate with eigenvalue $\hbar/2$ at all times. In other words, by rotating the observable $\hat{\psi}_{3}$ in synchrony with the precession of the spin (in fact, at twice the angular speed), one can continuously extract the same measurement result without disrupting the reachability domain.
In contrast, if we do not adjust the observable to follow the spin’s precession and simply measure $\hat{\psi}_{3}$, the system evolves into different superpositions over time, and after a full cycle $t=\frac{2\pi}{\tilde{B}}$, the state traverses the reachability domain $S^{1}$.
Similarly, if we instead choose $\hat{\psi}_{2}$ as the observable, the state is projected into an eigenstate of $\hat{\psi}_{2}$ at $t=0$, and its subsequent time evolution generates a different reachability domain $S^{\prime}{}^{1}$, which is another great circle on the Bloch sphere rotated around the same axis $\vec{n}$.
The orientation—or inclination—of this reachability domain with respect to the Bloch sphere is determined by the choice of observable, and thus reflects the structure of measurability imposed by the observer.
3.5.1 Interpretation: Reachability Domains as Constructed Spaces
In summary, while the precessional motion of spin on the Bloch sphere is well known as a unitary trajectory in the conventional picture, our approach interprets this trajectory not merely as an abstract path in state space, but as a dynamically generated reachability domain—a spatial structure that becomes accessible through observational acts. Unlike static geometric backgrounds, this domain reflects the concrete interface between system and observer: it is defined by the interplay between time evolution and the chosen observable. By synchronizing the observable with the system’s dynamics, one constructs an effectively static spatial frame, whereas fixing the observable yields a dynamically expanding set of accessible outcomes. Thus, space in this framework is not presupposed but emerges as a context-dependent construct. This perspective offers a constructive alternative to classical realism and aligns naturally with the contextual and noncommutative features of quantum theory, suggesting a viable foundation for quantum-geometric approaches to space.
4 Conclusion and Discussion
In this paper, we have analyzed and reconstructed the concept of space through the lens of physical frameworks such as classical and quantum systems. Our central claim has been that dynamical evolution and observational acts together serve as fundamental building blocks in the formation of spatial structure.
This perspective reveals a deep and intricate connection between the concept of space in mathematics and that in physics. Rather than treating space as a fixed, pre-given background, we have shown that it may emerge from the interplay between system dynamics and the observer’s mode of engagement. This approach invites a rethinking of mathematical foundations in a way that is more aligned with the operational language of physics—a reconstruction of mathematics that is, in some sense, physically meaningful.
However, such a shift also highlights the need for a reconstruction of mathematics itself, one that moves beyond the static formulations characteristic of classical mathematics prior to the 20th century. Developing a framework in which space is not presupposed but dynamically and observationally constructed remains a major open challenge for future work.
Acknowledgments
I would like to express my gratitude to Associate Professor Shiro Komata and Dr. Tsukasa Yumibayashi for reading the manuscript and providing valuable references. I am also indebted to Professor Emeritus Akio Sugamoto for his insightful comments on discrete and concrete models. Additionally, I thank the members of the Mathematical Physics Seminar for their numerous constructive comments on the earlier version of this work, originally published as [3]. These suggestions have been reflected throughout the present paper. I am grateful to Kazuya Mitsutani for his helpful comments on the Introduction.
Appendix A The GNS Construction and the Constructive Understanding of Space
In this appendix, we consider the formulation of quantum theory as a dynamical system, focusing on the construction of a corresponding Hilbert space given a state and a dynamical algebra. This construction is known as the GNS (Gelfand–Naimark–Segal) representation.
From the perspective of noncommutative geometry, the idea that space is nothing but algebraic structure has been strongly advocated—see Connes[5] or a foundational treatment. In particular, our position in this paper—that spatial structure emerges from the commutative part of observables—can be seen as a noncommutative generalization of the classical Gelfand–Naimark duality for commutative $C^{*}$-algebras. An accessible introduction to this viewpoint is provided in [6].
It is crucial to emphasize here that the definition of space depends on the observer’s choice of observable. The resulting spatial structure is not absolute, but contingent on the measurement context imposed by the observer.
Quantum theory, in this sense, can be viewed as a dynamical system in terms of observables $\hat{x}$. The essential shift from classical to quantum theory lies in the fact that these observables are no longer assumed to be mutually commutative.
To describe the dynamics of a quantum system, we begin by introducing a *-algebra as the algebraic structure of observables. For a rigorous definition and detailed proofs, see Bratteli and Robinson [4]. In Dirac’s terminology, this corresponds to moving from a dynamical system in the world of c-numbers (classical numbers) to one defined by q-numbers (quantum operators). Put more intuitively, this is equivalent to representing observables as complex matrices.In finite-dimensional systems, such as spin systems, this identification is exact: observables are literally represented as complex matrices acting on a Hilbert space.
Let us denote the collection of all observables by $\hat{X}$, with $\hat{x}\in\hat{X}$. A *-algebra is an algebra equipped with an involutive operation, denoted $\dagger$ (or * in mathematical notation), which satisfies certain algebraic properties. Specifically, applying $\dagger$ twice returns the original element:
$\left(\hat{x}^{\dagger}\right)^{\dagger}=\hat{x}$. (A.1)
In finite-dimensional systems, this operation corresponds to complex conjugate transpose—that is, Hermitian adjoint—of matrices:
$\hat{x}\mapsto\hat{x}^{\dagger}.$ (A.2)
The adjoint operation satisfies the following properties:
  • For linear combinations:
    $\left(a\hat{x}_{1}+b\hat{x}_{2}\right)^{\dagger}=a^{*}\hat{x}_{1}^{\dagger}+b^{*}\hat{x}_{2}^{\dagger},$ (A.3)
    where $a^{*}$ is denotes the complex conjugate of $a$.
  • For products:
    $\left(\hat{x}_{1}\hat{x}_{2}\right)^{\dagger}=\hat{x}_{2}^{\dagger}\hat{x}_{1}^{\dagger}.$ (A.4)
An observable $\hat{x}$ satisfying $\hat{x}=\hat{x}^{\dagger}$ is called self-adjoint, and corresponds to a physical quantity that can be measured.
The time evolution of an observable in quantum mechanics is written as:
$\hat{x}(t)=U(\hat{x}_{0},t),\ \mathrm{with\ }U(\hat{x}_{0},0)=\hat{x}_{0}.$ (A.5)
In the case of a Hamiltonian dynamical system, this evolution is typically given by the Heisenberg picture:
$\hat{x}(t)=e^{it\hat{H}}\hat{x}_{0}e^{-it\hat{H}}$, (A.6)
where $\hat{H}$ is a self-adjoint Hamiltonian operator generating the unitary time evolution.
The expectation value of an observable $\hat{x}$ is encoded by a state functional $\omega$, defined as:
$\omega(\hat{x})=\langle\hat{x}\rangle$. (A.7)
We require that $\omega(1)=1$, ensuring normalization. However, since the observable $\hat{x}$ may not commute with others, this expectation value cannot always be interpreted directly as a classical measure. To connect such functionals to Hilbert space structure, we invoke the GNS (Gelfand–Naimark–Segal) representation.
The GNS construction is a systematic procedure that, given a *-algebra $\mathcal{A}$ and a state $\omega$, produces a Hilbert space representation in which $\omega$ is realized as an inner product state.
Specifically, consider products of the form $\hat{x}^{\dagger}\hat{y}$. We interpret this as the inner product between two vectors:
$\omega(\hat{x}^{\dagger}\hat{y})=\langle\hat{x}|\hat{y}\rangle$, (A.8)
[…]
However, this construction faces a subtle issue: it is possible that
$\langle\hat{x}|\hat{x}\rangle=\omega(\hat{x}^{\dagger}\hat{x})=0,$ (A.9)
even though $\hat{x}\neq 0$ in the algebra. To resolve this, we define a subspace $\mathcal{N}_{\omega}\subset\mathcal{A}$, consisting of all null vectors:
$\mathcal{N}_{\omega}=\{\hat{x}_{\mathrm{null}}\in\mathcal{A}|\omega(\hat{x}^{\dagger}\hat{x})=0\}.$ (A.10)
We then define an equivalence relation:
$|\hat{x}\rangle\sim|\hat{x}+\hat{x}_{\mathrm{null}}\rangle$ (A.11)
and construct equivalence classes
$|[\hat{x}]_{\omega}\rangle\in\mathcal{A}/\mathcal{N}_{\omega}.$ (A.12)
A distinguished unit vector is given by the class corresponding to the identity:
$|\Omega_{\omega}\rangle=|[1]_{\omega}\rangle\sim|[1+\hat{x}_{\mathrm{null}}]_{\omega}\rangle,$ (A.13)
which satisfies the normalization condition:
$\langle\Omega_{\omega}|\Omega_{\omega}\rangle=\omega(1)=1.$ (A.14)
Mathematically, the inner product space constructed from the equivalence classes can be completed to form a Hilbert space, denoted $\mathcal{H}_{\omega}$.
We now define the action of an observable $\hat{x}\in\mathcal{A}$ on this Hilbert space by left multiplication:
$\pi_{\omega}(\hat{x})|[\hat{y}]\rangle=|[\hat{x}\hat{y}]_{\omega}\rangle.$ (A.15)
This definition turns $\pi_{\omega}$ into a *-representation of algebra $\mathcal{A}$ on the Hilbert space $\mathcal{H}_{\omega}$. In particular, we recover the equivalence class of $\hat{x}$ itself as:
$|[\hat{x}]_{\omega}\rangle=\pi_{\omega}(\hat{x})|[\hat{1}]_{\omega}\rangle.$ (A.16)
Consequently, the expectation value of an observable in the state $\omega$ can be written as:
$\omega(\hat{x})=\langle\hat{x}\rangle=\langle\Omega_{\omega}|\pi_{\omega}(\hat{x})|\Omega_{\omega}\rangle=\langle\Omega_{\omega}|[\hat{x}]_{\omega}\rangle,$ (A.17)
where $|\Omega_{\omega}\rangle=|[1]_{\omega}\rangle$ is the distinguished cyclic vector.
The pair $(\mathcal{H}_{\omega},\pi_{\omega})$ is called the GNS representation of the state $\omega$. Since the entire Hilbert space is generated by acting on the cyclic vector $|\Omega_{\omega}\rangle$, this representation is also known as a cyclic representation.
Let us now consider a simple example of a *-algebra in quantum mechanics: the quantum harmonic oscillator. This system is defined using self-adjoint operators $\hat{x}$ and $\hat{p}$ satisfying the canonical commutation relation:
$\hat{x},\hat{p}]=i\hbar.$ (A.18)
We introduce the creation and annihilation operators:
[…]
which satisfy the fundamental commutation relation:
$\hat{a},\hat{a}^{\dagger}]=1$ (A.20)
Any element $\hat{A}\in\hat{X}$ in *-algebra of observables can be written as a (possibly infinite) linear combination:
$\hat{A}=\sum_{n,m=0}^{\infty}c_{n,m}\hat{a}^{\dagger n}\hat{a}^{m}$ (A.21)
Now consider a state $\omega_{0}$ such that:
$\omega_{0}(\hat{a}^{\dagger}\hat{a})=0$ (A.22)
This implies:
$\langle\hat{a}|\hat{a}\rangle=0,$ (A.23)
so the equivalence class of $\hat{a}$ is null: $|\hat{a}\rangle\sim 0$. Then, the cyclic vector becomes:
$|\Omega_{\omega_{0}}\rangle=|[1]_{\omega_{0}}\rangle\sim|[1+\hat{a}]_{\omega_{0}}\rangle,$ (A.24)
and we have:
$\pi_{\omega_{0}}(\hat{a})|\Omega_{\omega_{0}}\rangle=|[\hat{a}]_{\omega_{0}}\rangle=0.$ (A.25)
More generally, we can construct a basis of states as:
$|n\rangle=\frac{1}{\sqrt{n!}}(\pi_{\omega_{0}}(\hat{a}^{\dagger}))^{n}|\Omega_{\omega_{0}}\rangle$ (A.26)
In this way, all physical states are generated by acting repeatedly with the creation operator on the cyclic vector, which physically corresponds to the vacuum state.
Thus, even in systems involving noncommutative observables, the GNS construction yields a Hilbert space $\mathcal{H}_{\omega}$ determined by the state $\omega$. That is, space emerges as a structure constructed from the observer’s perspective, depending on the algebraic properties of the observables and the choice of state.
Appendix B Commutativity and the Emergence of Classical Space: A Reconsideration via Quantum Observation
In this appendix, we reconsider the distinction between measurable spaces in classical theory and those in quantum theory. We argue that the essential difference lies solely in the noncommutativity of observables: the structure of measurable space is not fixed once and for all, but rather emerges anew each time an observer selects a specific observable.
From this perspective, every act of observation brings forth a new measure space—a classical (phase) space tailored to the set of compatible observables. This leads to a context-dependent view of space: space is not absolute, but arises from the particular measurement context determined by the observer’s choices.
However, such contextuality also raises the possibility that these constructed measure structures may not always be consistently compatible with one another. For further discussion of this issue, including the BKS (Bell–Kochen–Specker) theorem and the KCBS inequality, see the analysis by Kitano (2018) [7].
[…]
To state this viewpoint succinctly:
Proposition.
Observation is the act of constructing space.
In the following discussion, we adopt a formal approach without assuming boundedness of the *-algebra under consideration. For completeness, we note that a *-algebra with bounded elements is called a $C^{*}$-algebra, and that among these, those with particularly favorable properties are referred to as von Neumann algebras.
Since we do not assume boundedness in this paper, our arguments are not strictly confined to $C^{*}$-alebras. Rather, they presuppose the spectral decomposition over an appropriately extended algebraic structure, such as a (generalized) $W^{*}$-algebra. Thus, what follows should be understood as a schematic or formal treatment, intended to capture the essential structure of how classical space emerges in quantum theory through observable selection.
From the previous discussion, a quantum system is characterized by a *-algebra $\mathcal{A}$ of observables and a state $\omega$, where the observables $\hat{x}\in\mathcal{A}$ are self-adjoint. However, in this abstract algebraic formulation, the quantum system is not yet explicitly connected to a classical measure space.
In the classical case, the expectation value of an observable $\hat{x}$ is written as:
$\omega(\hat{x})=\langle x\rangle_{\rho}=\int xd\mu(x)=\int x\rho(x)dx$ (B.1)
where $\rho(x)$ is a classical probability density and $\mu$ is the associated measure.
In quantum theory, however, the expectation value takes the form:
[…]
Here, the object $\hat{\rho}_{\omega}=|\Omega_{\omega}\rangle\langle\Omega_{\omega}|$ is a density matrix, which serves as the quantum analogue of the classical probability density $\rho$. It encodes all the statistical information of the quantum state and allows us to express expectations in terms of operator traces rather than integrals over a predefined space.
Let us now examine how objects such as $x$, the underlying set, and $(X,\,\mathcal{B})$, the measurable space, appear within quantum theory. This requires a somewhat formal discussion, but it can be made precise by expressing the relevant quantities through the spectral decomposition of an observable $\hat{x}$ in terms of its eigenstates.
The crucial point is that this construction depends on the choice of observable $\hat{x}$. By projecting the density matrix onto the eigenbasis of $\hat{x}$, we extract those quantities that are measurable in the context of that observable. Specifically, the expectation value becomes:
$\omega(\hat{x})=\mathrm{Tr}\hat{\rho}\hat{x}=\sum_{x\in B(\hat{x})}\langle x|\hat{\rho}\hat{x}|x\rangle=\sum_{x\in B(\hat{x})}x\langle x|\hat{\rho}|x\rangle,$ (B.3)
[…]
In this framework, we see that the diagonal elements of the density matrix in the eigenbasis of $\hat{x}$ define a measure $\mu(\hat{x}$) , and thus construct a measurable space:
$).(X(\hat{x}),B(\hat{x}),\mu(\hat{x}$)). (B.5)
This reveals a profound difference between classical and quantum theories: namely, the role of noncommutativity. In classical mechanics, all observables commute and share a common measurable space. In quantum theory, however, each observable (or set of commuting observables) generates its own context-specific measure space—there is no globally fixed underlying space independent of observational choice.
For example, in classical mechanics, the expectation values of position $x$ and and momentum $p$ are given by:
$\omega_{\rho}(\hat{x})=\int x\rho(x,p)dxdp=\int xd\mu(x,p),$ (B.6)
$\omega_{\rho}(\hat{p})=\int p\rho(x,p)dxdp=\int pd\mu(x,p),$ (B.7)
where $(x,p)$ are points in phase space, distinguished by an appropriate Borel $\sigma$-algebra $\mathcal{B}$, and where the underlying space—here, the phase space—is denoted by $\Omega$. In this way, we obtain a classical measure space $(\Omega,\ \mathcal{B},\ \mu)$.
From the viewpoint of quantum theory, this corresponds to the assumption that position and momentum commute:
$\hat{x},\hat{p}]=0.$ (B.8)
That is, classical mechanics can be regarded as a special case of *commutative -algebra, in which all observables can be simultaneously diagonalized, and hence have a common system of joint eigenstates.
The expectation value in quantum theory then takes the form:
$\omega_{\rho}(\hat{x})=\mathrm{Tr}\hat{\rho}\hat{x}=\sum_{(x,p)\in B}\langle x,p|\hat{\rho}\hat{x}|x,p\rangle=\sum_{(x,p)\in B}x\rho((x,p),(x,p))$ (B.9)
[…]
Thus, when all observables under consideration are mutually commuting, we may select a common eigenbasis and thereby construct a single fixed measurable space:
$).(\Omega,B,\mu)=(X(\hat{x},\hat{p}),B(\hat{x},\hat{p}),\mu(\hat{x},\hat{p}$)). (B.11)
This is precisely how classical mechanics arises from within quantum theory, and it is the content of the Gelfand–Naimark theorem for commutative *-algebras: any commutative *-algebra is isomorphic to an algebra of functions over a classical topological (or measurable) space.
From the perspective of (phase) space, quantum theory inverts the classical logic: it begins not with a pre-given space, but with a *-algebra $\mathcal{A}$ of operators (i.e., q-numbers) and a state $\omega$ or density operator $\hat{\rho}$ that defines expectation values. In other words, quantum theory is specified by a pair $(\mathcal{A},\omega$), where structure and dynamics are encoded algebraically.
Given this, the observer selects a self-adjoint observable $\hat{a}\in\mathcal{A}$ (i.e., $\hat{a}=\hat{a}^{\dagger})$ to measure. Each such choice induces a corresponding measure space:
$),(X(\hat{a}),B(\hat{a}),\mu(\hat{a}$)), (B.12)
which may be interpreted as a classical world emergent from the context defined by $\hat{a}$.
This viewpoint—that the structure of space emerges from the observer’s act of measurement—resonates with Rovelli’s relational quantum mechanics [9, 10, 11], where space is not treated as a fixed background but as a web of relations defined through interactions. In this sense, our approach represents a serious reevaluation of the quantum worldview, combining the Heisenberg picture with the Born rule and grounding spatial structure in operational choices.
This leads us naturally to further questions:
Problem.
What kinds of sets $\{\hat{a}_{i}\}_{i=1}^{N}$ of observables can be simultaneously measured in quantum theory? How do these sets transform under time evolution?
The structure of such commutative subalgebras within a noncommutative algebra is a deep subject in quantum foundations. The pioneering results of Kochen and Specker [19] showed that no global assignment of values to all observables is possible in general, due to contextuality. For a detailed philosophical and information-theoretic discussion, we refer the reader to Bub’s analysis [20], where the logical structure and classical reality conditions of quantum theory are thoroughly examined.